# Early Warning Signals for Bifurcations Embedded in High Dimensions

### v2

@@ Abstract
Recent work has highlighted the utility of methods for early warning signal detection in dynamic systems approaching critical tipping thresholds. Often these tipping points resemble local bifurcations, whose low dimensional dynamics can play out on a manifold embedded in a much higher dimensional state space. In many cases of practical relevance, the form of this embedding is poorly understood or entirely unknown. This paper explores how measurement of the critical phenomena that generically precede such bifurcations can be used to make inferences about the properties of their embeddings, and, conversely, how prior knowledge about the mechanism of bifurcation can robustify predictions of an oncoming tipping event. These modes of analysis are first demonstrated on a simple fluid flow system undergoing a Hopf bifurcation. The same approach is then applied to data associated with the West African monsoon shift, with results corroborated by existing models of the same system. This example highlights the effectiveness of the methodology even when applied to complex climate data, and demonstrates how a well-resolved spatial structure associated with the onset of atmospheric instability can be inferred purely from time series measurements.
@@ Introduction
Many dynamical systems contain critical thresholds which act as “tipping points,” parameter boundaries whose traversal induces an abrupt and potentially irreversible shift in the state. These critical transition events pose considerable concern in climatic, ecological, economic, and other domains in which a sudden, dramatic reconfiguration could have grave consequences for human welfare. Increasing attention has recently been devoted to understanding how early warning signals (EWS) associated with an oncoming tipping point may manifest in a system’s dynamics before it reaches criticality.
@@
This paper aims to address the question of how EWS analysis of complex systems might be better informed by explicit attention to the embedding of the bifurcation manifold in their state spaces. In particular, we investigate how inferences can be made in either direction between a hypothesis of tipping mechanism and the observation of critical phenomena. In cases where properties of a potential bifurcation can be hypothesized, we discuss how these hypotheses can be used to refine and robustify the EWS detection process. Conversely, in the absence of prior knowledge about the tipping scenario, we demonstrate how EWS methodologies can be employed to make indirect observations of the bifurcation embedding. We focus our discussion on local bifurcations taking place in high dimensional spaces, with particular attention to systems where spatiotemporally resolved measurements are made on one or more dynamical variables.
These cases are highly relevant to studies of climate tipping phenomena, where extensive geospatial data has been made available by the modern proliferation of remote sensing technology. Tipping points in the climate have been the subject of a great deal of recent interest amid a growing consensus that their potential abruptness and unpredictability places them among the gravest threats faced in a warming world [1, 2, 3, 4]. A central goal of this research is the identification of “tipping elements”—components of the Earth system which may be prone to a runaway feedback mechanism when forced past some critical threshold [5]. Using the example of an abrupt shift observed in the annual West African monsoon system, we demonstrate how our method for inferring the structure of an embedded bifurcation can aid in this task. We employ EWS analysis to obtain a well-resolved spatial representation of the support of the critical dynamics directly from observation data, validating our results against those of a more conventional modeling approach.
@@
The basis of EWS detection techniques is the observation that many types of bifurcation are guaranteed to share certain characteristics, which can produce detectable signatures [6, 7]. All local bifurcations in continuous time, for example, occur when an eigenvalue of the Jacobian matrix crosses the imaginary axis. A system in a stable equilibrium approaching a local bifurcation will find its basin of attraction growing increasingly shallow, resulting in reduced resilience and longer recovery times in response to small perturbations. This phenomenon of “critical slowing down” (CSD) can be observed statistically in time series measurements, typically by examining trends in variance or autocorrelation [8].
@@
Bifurcations are taxonomized according to the different types of changes that the topological structure of a phase portrait can undergo. Any class of local bifurcation can be described by a normal form, a differential equation which minimally reproduces the bifurcation in as few dimensions as possible [9]. All instances of a given type of local bifurcation are guaranteed to be topologically equivalent to the normal form near to the equilibrium. Crucially, this holds true even for high-dimensional systems: the bifurcation dynamics represented by the normal form are reproduced on a low-dimensional manifold embedded in the state space. Consider a system with dynamics given by
$\frac{d\mathbf{x}}{dt}=\mathbf{f}\left(\mathbf{x},\alpha\right)$ (1)
where $\mathbf{f}:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{n}$ is some nonlinear function, $\mathbf{x}$ is the $n$-dimensional state variable, and $\alpha$ is a bifurcation parameter which we take to be a real scalar. Given that the system undergoes some local bifurcation at $(\mathbf{x}_{c},\alpha_{c})$, there exists a projector function $\mathbf{g}:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{m}$ such that the dynamics close to the bifurcation point satisfy
$\frac{d\mathbf{g}(\mathbf{x},\alpha)}{dt}=N_{B}\left(\mathbf{g}(\mathbf{x},\alpha),\alpha\right)$ (2)
where $N_{B}:\mathbb{R}^{m}\times\mathbb{R}\to\mathbb{R}^{m}$ is the right-hand side of the normal form of the bifurcation (which occurs in $m\leq n$ dimensions).
In practical terms, critical transitions in complex systems are often only approximated by local bifurcations. A system with dynamics on highly disparate time scales, for example, might admit an approximate representation for its slow manifold which reduces to the form of a bifurcation after averaging over the fast-scale dynamics. In these cases, the bifurcation embedding is a manifold on which observed dynamics are approximated imperfectly by the bifurcation normal form.
@@
With respect to tipping phenomena in climate systems, the embedding (exact or approximate) of the bifurcation manifold within the state space is typically unknown. EWS detection methods are usually carried out using direct measurements of the system state under the assumption that the critical phenomena produced by the bifurcation dynamics in Eq. 2 will be reproduced in the full state dynamics of Eq. 1. This is, in many cases, a well-motivated assumption. Indications of an oncoming bifurcation are usually encoded in trends in variance, autocorrelation, skewness, and other statistical properties which are fairly robust to transformation: an increase in variance of $\mathbf{g}(\mathbf{x},\alpha)$ on the bifurcation manifold is very likely to correspond to an increase in variance of $\mathbf{x}$ for any reasonably smooth and well-behaved function $\mathbf{g}$. As a result, early warning for a tipping event can often be successfully measured even without any knowledge of the embedding $\mathbf{g}$. Nevertheless, any attempt to monitor EWS entirely agnostically of $\mathbf{g}$ carries some risk: if one unknowingly measures the system state along a direction orthogonal to the bifurcation manifold, the signs of oncoming tipping can be completely obscured. This phenomenon of “silent catastrophe” has been demonstrated in low-dimensional ecological models by Boerlijst et. al. (2013) [10], but it leads to potentially greater difficulty in high dimensional cases, where large subspaces of unknown structure may be devoid of EWS information.
@@
EWS Detection Methods
Results presented in this work are based on the application of an EWS detection model to time series data. A large variety of methodologies have been proposed for measuring signatures of oncoming critical transitions, diverse in both the types of transitions they are intended to forecast (local bifurcations, nonlocal bifurcations, phase transitions, etc.) [7, 11, 12] and in the analytical means by which they do so (temporal or spatial computation of autocorrelation, moments of distribution, spectral methods, etc.) [8, 13]. Additionally, recent work has shown the potential of machine learning models to improve on these traditional statistical approaches by improving on accuracy and offering greater insight into the nature of the oncoming transition [14, 15, 16]. In this paper we employ a neural model of our own construction, based primarily on the design of Bury et. al. (2021) [14]. The classifier takes as input a univariate time series of arbitrary length, and outputs a vector of probabilities assigned to each of the target classes: null (no upcoming bifurcation) or any of four types of local bifurcation (fold, Hopf, transcritical, and pitchfork). Details on the architecture and training of this model are presented in Appendix A.
We chose a deep learning approach for its classification accuracy and flexibility, but these methods are not without drawbacks. Neural networks are well known to be fairly opaque in their execution, sacrificing interpretability for performance. Furthermore, they tend to overstate their confidence in results for classification tasks, i.e. outputting class probabilities very close to 0% or 100% even when their demonstrated accuracy on validation tasks is considerably lower (though this can be mitigated somewhat through calibration methods such as temperature scaling) [17]. We emphasize that none of the analysis in this paper is specific to any one EWS detection method. All results we present are produced by our neural classifier model, but the same methodologies could be equally effectively carried out using any other approach (albeit subject to the strengths and weaknesses of the chosen technique).
@@
Figure 1: Schematic of the embedding of the 2D Hopf bifurcation dynamics in the fluid system. Left: vorticity plots of the full system state in its laminar equilibrium (top) and vortex shedding oscillation (bottom). Center: vorticity plots illustrating the spatial modes which form an approximate linear basis for the bifurcation manifold. Right: dynamics of the system projected into the 3D space spanned by these modes (black) overlaid on the 2D manifold on which the Hopf dynamics unfold (blue)
As a demonstrative example, we present results for the application of EWS methods to a simulated fluid flow system undergoing a bifurcation. A fluid in two dimensions flows at uniform velocity towards a circular obstacle. By numerical simulation of the 2D Navier-Stokes equations, two distinct regimes can be observed: low flow rates produce a steady, laminar wake behind the obstacle, whereas high flow rates lead to a oscillatory solution known as a von Kármán vortex street, in which eddies of alternating vorticity are periodically shed past the obstacle. The abrupt transition of the system from a fixed equilibrium to a stable periodic orbit due to smooth variation of the flow rate bears all the hallmarks of a Hopf bifurcation, but the mechanism by which it occurs is not obvious: a Hopf bifurcation is produced by a cubic nonlinearity, whereas the Navier-Stokes equations governing the fluid system only contain terms up to second order. This apparent contradiction led to a lengthy debate among fluid dynamicists which was ultimately resolved by the development of a mean-field model which isolates a slow timescale manifold on which dynamics approximate the cubic term of the Hopf normal form [18, 19]. Details of this solution are presented in Appendix B.
@@
In the absence of prior knowledge about the bifurcation embedding, EWS detection is generally carried out using data from direct measurements on the state space. Figure 2 shows results for the neural EWS model applied to the simulated fluid flow system approaching its Hopf bifurcation. EWS probabilities are computed separately for the observed dynamics at each spatial coordinate on a sliding time window leading up to the critical point. Results averaged over the spatial dimensions (bottom plot) show that the model has performed as expected: probability for the Hopf class remains low throughout most of the laminar flow regime and increases significantly as the system nears its vortex shedding transition. The bifurcation is successfully predicted using measurements of both horizontal and vertical components of the flow, with the latter slightly outperforming the former in its forecast horizon. This discrepancy may result from a difference in signal-to-noise ratio; dominant mean flow is in the horizontal direction, so the comparatively small perturbative excursions in which much of the EWS information is encoded may be easier to discern along the vertical axis.
Figure 2: Top: envelope of vortex shedding cycle in the 3D mean field model, projected onto the full state space of horizontal flow (left) and vertical flow (right). Middle: snapshot of the EWS detection model’s spatially-resolved output probability of an oncoming Hopf bifurcation based on measurements of horizontal flow rate (left) and vertical flow rate (right). Bottom: spatially averaged model output over time as the system approaches its bifurcation (denoted by the black vertical line). The times of the middle snapshots are represented by the dashed vertical lines
@@
The second reason for possible disagreement is based on the fact that the principal mechanism by which early warning signals manifest in dynamics is through the system’s response to perturbations. Consequently, the ability to detect critical phenomena from measurement data depends not only on the system’s endogenous dynamics as it nears bifurcation, but also on the magnitude and spectral properties of any forcing it is experiencing. In the case of this fluid flow system a very low level of white noise is applied at each step of the simulation, but perturbations to the bifurcation mechanics are primarily contributed by dynamics outside the subspace spanned by $(\mathbf{u}_{w},\mathbf{u}_{x},\mathbf{u}_{y})$, which can be considered exogenous to the bifurcation. These dynamics, resulting from the discrepancies between the full Navier-Stokes evolution and its mean field approximation, have a spatial character of their own. This spatial heterogeneity is likely to cause the EWS model’s sensitivity to vary by location in a manner that would not be captured by the mean field predictions plotted in Fig. 2. This phenomenon represents a potential impediment to inferring the structure of a bifurcation embedding purely through EWS analysis. It is worth noting, however, that the plots in Fig. 2 are snapshots of model output well in advance of the onset of the bifurcation. Closer to criticality, the yellow portions of these plots expand considerably and fill nearly the entire region to the right of the cylinder. Even if sensitivity has been adversely impacted by nonuniform perturbative forcing, the model ultimately succeeds in correctly classifying its input as warning signals become more pronounced approaching the bifurcation.
@@
We have demonstrated how EWS analysis can be employed to draw inferences about an unknown embedding of a bifurcation in some measurement space, but the same relationship can also be usefully applied in the other direction: if one can leverage prior knowledge to make certain assumptions about how a bifurcation will manifest in some system of interest, those can be used to robustify EWS detection to predict the critical transition. As an illustrative example, we again consider the Hopf bifurcation of the same 2D fluid system and assume that the mean field solution is not known. Even without knowledge of the “proper” bifurcation basis, one can reasonably expect that the spatial embedding of the bifurcation will contain structures on the same order of magnitude as the characteristic length scale of the system, i.e., the diameter of the cylinder. This offers an additional constraint that can be placed on EWS analysis of the dynamics: if probability of an oncoming critical transition is computed separately at each spatial coordinate, one can evaluate the likelihood of an incipient vortex shedding bifurcation by searching for a coherent rise in output probability across a region of the expected size.
@@
Figure 3: Top: Snapshot of the $x$ component (i.e. along the horizontal) of the turbulent channel flow system. Bottom: Snapshot of the spatially resolved output of the neural EWS model’s Hopf class
@@
Figure 4: Left: Time series of the neural EWS model’s Hopf class output computed at a randomly chosen spatial location in the vortex shedding system (top) and turbulent flow (bottom). Right: Time series predictions for the same systems computed by the spatially smoothed global method. For the vortex shedding system, the time of the true Hopf bifurcation is annotated by the dashed vertical lines
In Fig. 4 we present a simple example of how this difference might be operationalized for more robust EWS detection. Model output probability (for the Hopf class) over time for a single, arbitrary spatial location in each system is plotted on the left. Here we observe that the vortex shedding system yields a (true) positive classification shortly before the Hopf bifurcation, but the turbulent flow system also produces (effectively false) positive classifications at various times through its duration. To improve these results, we implement a simple postprocessing step: model outputs are smoothed along the spatial dimensions using a Gaussian convolution filter with width determined by the system’s characteristic length scale (cylinder diameter and half channel width, respectively). This effectively obscures more localized excitations, allowing for an approximate global probability to be inferred from the maximum value of the resulting signal. These global probability metrics, plotted in Fig. 4 (right), produce the desired outcome: positive classifications for the turbulence case are entirely suppressed, while those for the vortex shedding case are preserved and even enhanced, arising further in advance of the bifurcation.
This methodology is expected to be quite adaptable. The example presented relies on an assertion of the expected spatial scale of the bifurcation, but a similar approach could be applied for an assumed characteristic time scale or for any other circumstance in which prior knowledge about the system can be used to constrain the search for an unknown embedding a high-dimensional measurement space. EWS detection methods have a documented proclivity for erroneous positive classifications [25, 26], so any means of reducing such errors may constitute a considerable step forward for practical applications (although it should be noted that our method does not directly counter the sources of systematic error addressed in those references).
@@
Figure 5: Gradient of the neural model’s fold class output (average daily rate of change) between June 15 and July 8 (mean over all years) for measurements of the meridional component of neutral wind at 10 m above the surface of the Earth. The east Atlantic marine ITCZ and Guinean coast averaging regions used by Cook (2015) are labeled for reference
We use this result as a point of comparison for the spatial embedding inferred by the EWS methodology described in the previous section. Climate data is obtained from the ERA5 global reanalysis for the years 1940 to 2022 [35, 36]. Hourly data is downsampled to daily means to remove day-night oscillations. Following the reasoning of Cook, we focus our analysis on the meridional (north-south) component of wind: the dominant, approximately geostrophic flow in the (spatial and temporal) region of interest is primarily in the zonal (east-west) direction, so any critical phenomena produced by the proposed mechanism of inertial instability are expected to manifest in the system’s response to meridional perturbations. As with the vortex shedding system, model predictions are computed separately for each spatial location on a sliding time window (of length 60 days). The observed dynamics of the ERA5 data are of course the product of the many overlapping mechanisms of diverse spatial and temporal scale which drive the evolution of the Earth’s climate, any of which might produce signatures associated with critical transitions. In order to help isolate any signal connected with the dynamics connected to the monsoon shift phenomenon, we examine the gradient of model output averaged over a period of about 3 weeks prior to the expected onset time (roughly July 8, although it varies from year to year).

### v3

@@ Abstract
Recent work has highlighted the utility of methods for early warning signal detection in dynamic systems approaching critical tipping thresholds. Often these tipping points resemble local bifurcations, whose low dimensional dynamics can play out on a manifold embedded in a much higher dimensional state space. In many cases of practical relevance, the form of this embedding is poorly understood or entirely unknown. This paper explores how measurement of the critical phenomena that generically precede such bifurcations can be used to make inferences about some properties of their embeddings, and, conversely, how prior knowledge about the mechanism of bifurcation can robustify predictions of an oncoming tipping event. These modes of analysis are first demonstrated on a simple fluid flow system undergoing a Hopf bifurcation. The same approach is then applied to data associated with the West African monsoon shift, with results corroborated by existing models of the same system. This example highlights the effectiveness of the methodology even when applied to complex climate data, and demonstrates how a well-resolved spatial structure associated with the onset of atmospheric instability can be inferred purely from time series measurements.
@@ Introduction
Many dynamical systems contain critical thresholds which act as “tipping points," parameter boundaries whose traversal induces an abrupt and potentially irreversible shift in the state. These critical transition events pose considerable concern in climatic, ecological, economic, and other domains in which a sudden, dramatic reconfiguration could have grave consequences for human welfare. Increasing attention has recently been devoted to understanding how early warning signals (EWS) associated with an oncoming tipping point may manifest in a system’s dynamics before it reaches criticality.
@@
This paper aims to address the question of how EWS analysis of complex systems might be better informed by explicit attention to the embedding of a bifurcation in their state spaces. In particular, we investigate how inferences can be made in either direction between a hypothesis of tipping mechanism and the observation of critical phenomena. In cases where properties of a potential bifurcation can be hypothesized, we discuss how these hypotheses can be used to refine and robustify the EWS detection process (in the spatial aggregation step of the depiction in Box 1). Conversely, in the absence of prior knowledge about the tipping scenario, we demonstrate how EWS methodologies can be employed to make indirect observations of the support of the bifurcation embedding (bottom right panel of Box 1). Although previous EWS studies have in many cases made implicit use of assumed properties of the bifurcation in question (evidenced by their choice of measurement variables, sampling resolution, etc.), these hypotheses are rarely, if ever, contextualized in terms of their relationship to a bifurcation embedding. Likewise, the possibility of inferring embedding properties of a bifurcation directly from observed critical phenomena has gone largely unexplored in existing EWS literature.
We focus this study on local bifurcations taking place in high dimensional spaces, with particular attention to systems where spatiotemporally resolved measurements are made on one or more dynamical variables. The methods we introduce are entirely agnostic of the EWS detection algorithm employed. We present results obtained using a neural classifier model, as discussed in the Methods section, but the same analysis could equally be conducted using any other detection technique (albeit subject to the strengths and weaknesses of the chosen approach).
These cases are highly relevant to studies of climate tipping phenomena, where extensive geospatial data has been made available by the modern proliferation of remote sensing technology. Tipping points in the climate have been the subject of a great deal of recent interest amid a growing consensus that their potential abruptness and unpredictability places them among the gravest threats faced in a warming world [1, 2, 3, 4]. A central goal of this research is the identification of “tipping elements"—components of the Earth system which may be prone to a runaway feedback mechanism when forced past some critical threshold [5]. Using the example of an abrupt shift observed in the annual West African monsoon system, we demonstrate how our method for inferring the structure of an embedded bifurcation can aid in this task. We employ EWS analysis to obtain a well-resolved spatial representation of the support of the critical dynamics directly from observation data, validating our results against those of a more conventional modeling approach.
[t] [Uncaptioned image] Box 1: Conceptual schematic of embedded-bifurcation EWS analysis
@@
The basis of EWS detection techniques is the observation that many types of bifurcation are guaranteed to share certain characteristics, which can produce detectable signatures [6, 7]. All local bifurcations in continuous time, for example, occur when an eigenvalue of the Jacobian matrix crosses the imaginary axis. A system in a stable equilibrium approaching a local bifurcation will find its basin of attraction growing increasingly shallow, resulting in reduced resilience and longer recovery times in response to small perturbations. This phenomenon of “critical slowing down" (CSD) can be observed statistically in time series measurements, typically by examining trends in variance or autocorrelation [8].
@@
Bifurcations are taxonomized according to the different types of changes that the topological structure of a phase portrait can undergo. Any class of local bifurcation can be described by a normal form, a differential equation which minimally reproduces the bifurcation in as few dimensions as possible [9]. All instances of a given type of local bifurcation are guaranteed to be topologically equivalent to the normal form near to the equilibrium. Crucially, this holds true even for high-dimensional systems: the bifurcation dynamics represented by the normal form are reproduced on a low-dimensional manifold embedded in the state space.
In practical terms, critical transitions in complex systems are often only approximated by local bifurcations. A system with dynamics on highly disparate time scales, for example, might admit an approximate representation for its slow manifold which reduces to the form of a bifurcation after averaging over the fast-scale dynamics. In these cases, the bifurcation embedding is a manifold on which observed dynamics are approximated imperfectly by the bifurcation normal form. It should also be noted that although CSD phenomena are produced by all types of local bifurcations, not all of these events are accompanied by the discontinuous, irreversible changes most commonly associated with the term "tipping point." Distinguishing between signatures of oncoming bifurcations which will or will not produce an abrupt equilibrium shift remains an open problem in EWS detection research. Some methods have shown promise, including neural time series classification models such as the one used in this study, but further investigation is needed [10, 11]. For the purposes of this paper we primarily concern ourselves with the detection of any bifurcation transition, continuous or otherwise. In practical terms, even those bifurcations which do not induce sudden and irreversible regime shifts may nonetheless lead to significant qualitative changes in a system’s behavior on a relatively short time scale. While differentiating between the two remains an important goal, we emphasize that the results presented here do not offer any improvements on existing techniques to do so.
@@
With respect to climate systems, the embedding (exact or approximate) of a bifurcation manifold within the state space is typically unknown. EWS detection methods are usually carried out using direct measurements of the system state under the assumption that the critical phenomena produced by the low-dimension bifurcation dynamics will be reproduced in the full state dynamics. This is, in many cases, a well-motivated assumption. Indications of an oncoming bifurcation are usually encoded in trends in variance, autocorrelation, skewness, and other statistical properties which are fairly robust to transformation: an increase in variance along the critical axis of the low-dimensional bifurcation system is very likely to correspond to an increase in variance of a measurement made on the high-dimensional state space (for any reasonably smooth and well-behaved embedding of the bifurcation). As a result, early warning for a tipping event can often be successfully measured even without any knowledge of the embedding. Nevertheless, any attempt to monitor EWS entirely agnostically of this relationship carries some risk: if one unknowingly measures the system state along a direction orthogonal to the bifurcation manifold, the signs of oncoming tipping can be completely obscured. This phenomenon of “silent catastrophe" has been demonstrated in low-dimensional ecological models by Boerlijst et. al. (2013) [12], but it leads to potentially greater difficulty in high dimensional cases, where large subspaces of unknown structure may be devoid of EWS information.
@@
Results
@@
Figure 1: Schematic of the embedding of the 2D Hopf bifurcation dynamics in the fluid system. The high-dimensional state $\mathbf{x}$ is (approximately) reduced to the bifurcation form by the embedding function $\mathbf{g}$. Left: vorticity plots of the full system state in its laminar equilibrium (top) and vortex shedding oscillation (bottom). Center: vorticity plots illustrating the spatial modes which form an approximate linear basis for the bifurcation manifold. Right: dynamics of the system projected into the 3D space spanned by these modes (black) overlaid on the 2D manifold on which the Hopf dynamics unfold (blue).
As a demonstrative example, we present results for the application of EWS methods to a simulated fluid flow system undergoing a bifurcation. A fluid in two dimensions flows at uniform velocity towards a circular obstacle. By numerical simulation of the 2D Navier-Stokes equations, two distinct regimes can be observed: low flow rates produce a steady, laminar wake behind the obstacle, whereas high flow rates lead to a oscillatory solution known as a von Kármán vortex street, in which eddies of alternating vorticity are periodically shed past the obstacle. The abrupt transition of the system from a fixed equilibrium to a stable periodic orbit due to smooth variation of the flow rate bears all the hallmarks of a Hopf bifurcation, but the mechanism by which it occurs is not obvious: a Hopf bifurcation is produced by a cubic nonlinearity, whereas the Navier-Stokes equations governing the fluid system only contain terms up to second order. This apparent contradiction was resolved by the development of a mean-field model which isolates a slow timescale manifold on which dynamics approximate the cubic term of the Hopf normal form [13, 14]. Details of this solution are presented in the supporting materials.
@@
In the absence of prior knowledge about the bifurcation embedding, EWS detection is generally carried out using data from direct measurements on the state space. Figure 2 shows results for the neural EWS model applied to the simulated fluid flow system approaching its Hopf bifurcation. EWS probabilities are computed separately for the observed dynamics at each spatial coordinate on a sliding time window leading up to the critical point. It should be noted that this does forgo any information that might be encoded in spatial correlations, which has been shown to be useful for EWS detection in many spatiotemporal systems [15, 16]. We opt for this approach because treating each coordinate independently has the advantage of preserving spatial detail in the model output. Although we do not implement it here, a compromise between the two could be achieved by applying a spatiotemporal EWS model on a sliding spatial window over the data set. This would allow the model to make use of local spatial relationships, albeit at some cost to the spatial resolution of the results.
Results averaged over the spatial dimensions (bottom plot) show that the model has performed as expected: probability for the Hopf class remains low throughout most of the laminar flow regime and increases significantly as the system nears its vortex shedding transition. The bifurcation is successfully predicted using measurements of both horizontal and vertical components of the flow, with the latter slightly outperforming the former in its forecast horizon. This discrepancy may result from a difference in signal-to-noise ratio; dominant mean flow is in the horizontal direction, so the comparatively small perturbative excursions in which much of the EWS information is encoded may be easier to discern along the vertical axis.
Figure 2: Top: envelope of vortex shedding cycle in the 3D mean field model, projected onto the full state space of horizontal flow (left) and vertical flow (right). Middle: snapshot of the EWS detection model’s spatially-resolved output probability of an oncoming Hopf bifurcation based on measurements of horizontal flow rate (left) and vertical flow rate (right). Bottom: spatially averaged model output over time as the system approaches its bifurcation (denoted by the black vertical line). The times of the middle snapshots are represented by the dashed vertical lines.
@@
The second reason for possible disagreement is based on the fact that the principal mechanism by which early warning signals manifest in dynamics is through the system’s response to perturbations. Consequently, the ability to detect critical phenomena from measurement data depends not only on the system’s endogenous dynamics as it nears bifurcation, but also on the magnitude and spectral properties of any forcing it is experiencing. In the case of this fluid flow system a very low level of white noise is applied at each step of the simulation, but perturbations to the bifurcation mechanics are primarily contributed by dynamics outside the subspace spanned by $(\mathbf{u}_{w},\mathbf{u}_{x},\mathbf{u}_{y})$, which can be considered exogenous to the bifurcation. These dynamics, resulting from the discrepancies between the full Navier-Stokes evolution and its mean field approximation, have a spatial character of their own. This spatial heterogeneity is likely to cause the EWS model’s sensitivity to vary by location in a manner that would not be captured by the mean field predictions plotted in Fig. 2. This phenomenon represents a potential impediment to inferring the structure of a bifurcation embedding purely through EWS analysis. It is worth noting, however, that the plots in Fig. 2 are snapshots of model output well in advance of the onset of the bifurcation. Closer to criticality, the yellow portions of these plots expand considerably and fill nearly the entire region to the right of the cylinder. Even if sensitivity has been adversely impacted by nonuniform perturbative forcing, the model ultimately succeeds in correctly classifying its input as warning signals become more pronounced approaching the bifurcation. Furthermore, it is likely that a spatially convolutional EWS model could be trained on examples with heterogeneous forcing in order to mitigate this effect. As discussed previously, this would come at the expense of some spatial resolution in the output, but it is nonetheless an approach worth exploring in future work.
@@
We have demonstrated how EWS analysis can be employed to draw inferences about an unknown embedding of a bifurcation in some measurement space, but the same relationship can also be usefully applied in the other direction: if one can leverage prior knowledge to make certain assumptions about how a bifurcation will manifest in some system of interest, those can be used to robustify EWS detection to predict the critical transition. As an illustrative example, we again consider the Hopf bifurcation of the same 2D fluid system and assume that the mean field solution is not known. Even without knowledge of the “proper" bifurcation basis, one can reasonably expect that the spatial embedding of the bifurcation will contain structures on the same order of magnitude as the characteristic length scale of the system, i.e., the diameter of the cylinder. This offers an additional constraint that can be placed on EWS analysis of the dynamics: if probability of an oncoming critical transition is computed separately at each spatial coordinate, one can evaluate the likelihood of an incipient vortex shedding bifurcation by searching for a coherent rise in output probability across a region of the expected size.
@@
Figure 3: Top: Snapshot of the $x$ component (i.e. along the horizontal) of the turbulent channel flow system (arbitrary units). Bottom: Snapshot of the spatially resolved output of the neural EWS model’s Hopf class. The model’s spatially-averaged output probability at any given time is quite low, but there are many localized regions in which it assigns a high likelihood of an oncoming bifurcation.
@@
Figure 4: Left: Time series of the neural EWS model’s Hopf class output computed at a randomly chosen spatial location in the vortex shedding system (top) and turbulent flow (middle). Right: Time series predictions for the same systems computed by the spatially smoothed global method. For the vortex shedding system, the time of the true Hopf bifurcation is annotated by the dashed vertical lines. Bottom: ROC curves computed from an ensemble of randomly sampled spatial locations in both systems. In the smoothed case, this is carried out before computing the global maximum in order to allow for spatial sampling.
In Fig. 4 we present a simple example of how this difference might be operationalized for more robust EWS detection. Model output probability (for the Hopf class) over time for a single, arbitrary spatial location in each system is plotted on the left. Here we observe that the vortex shedding system yields a (true) positive classification shortly before the Hopf bifurcation, but the turbulent flow system also produces (effectively false) positive classifications at various times through its duration. To improve these results, we implement a simple postprocessing step: model outputs are smoothed along the spatial dimensions using a Gaussian convolution filter with width determined by the system’s characteristic length scale (cylinder diameter and half channel width, respectively). This effectively obscures more localized excitations, allowing for an approximate global probability to be inferred from the maximum value of the resulting signal. These global probability metrics, plotted in Fig. 4 (right), produce the desired outcome: positive classifications for the turbulence case are entirely suppressed, while those for the vortex shedding case are preserved and even enhanced, arising further in advance of the bifurcation. This improvement is further demonstrated in the bottom row of Fig. 4, where receiver operating characteristic (ROC) curves are produced from ensemble statistics over a large number of randomly sampled spatial locations in each system (in the right column, this sampling is carried out before the global maximum is computed). The area under the curve is observed to increase from 0.84 to 1.
This postprocessing methodology is expected to be quite adaptable. The example presented relies on an assertion of the expected spatial scale of the bifurcation, but a similar approach could be applied for an assumed characteristic time scale or for any other circumstance in which prior knowledge about the system can be used to constrain the search for an unknown embedding a high-dimensional measurement space. EWS detection methods have a documented proclivity for erroneous positive classifications [22, 23], so any means of reducing such errors may constitute a considerable step forward for practical applications (although it should be noted that our method does not directly counter the sources of systematic error addressed in those references).
@@
Figure 5: Gradient of the neural model’s fold class output (average daily rate of change) between June 15 and July 8 (mean over all years) for measurements of the meridional component of neutral wind at 10 m above the surface of the Earth. The detected onset of dynamic instability occurs most prominently just off the coast to the south and west of the mainland. The east Atlantic marine ITCZ and Guinean coast averaging regions used by Cook (2015) are labeled for reference.
We use this result as a point of comparison for the spatial embedding inferred by the EWS methodology described in the previous section. Climate data is obtained from the ERA5 global reanalysis for the years 1940 to 2022 [32, 33]. Hourly data is downsampled to daily means to remove day-night oscillations. Initial results make use of measurements of neutral wind at a height of 10m above the Earth’s surface, with further discussion of variable selection to follow. Neutral wind is computed based on surface stress to reflect air flow in the idealized case of neutral atmospheric stratification. Following the reasoning of Cook, we focus our analysis on the meridional (north-south) component of wind: the dominant, approximately geostrophic flow in the (spatial and temporal) region of interest is primarily in the zonal (east-west) direction, so any critical phenomena produced by the proposed mechanism of inertial instability are expected to manifest in the system’s response to meridional perturbations. As with the vortex shedding system, model predictions are computed separately for each spatial location on a sliding time window (of length 60 days). The observed dynamics of the ERA5 data are of course the product of the many overlapping mechanisms of diverse spatial and temporal scale which drive the evolution of the Earth’s climate, any of which might produce signatures associated with critical transitions. In order to help isolate any signal connected with the dynamics connected to the monsoon shift phenomenon, we examine the gradient of model output averaged over a period of about 3 weeks prior to the expected onset time (roughly July 8, although it varies from year to year).